# Heterogeneous Treatment Effects and

### v1

@@ Appendix C Directly/Indirectly-Affected versus Latent/Observed Outcomes
Appendix D Comparison to Mediation Analysis
In this section, we compare our framework connecting HTEs and mechanisms to mediation analysis. It is important to note that the two frameworks are built on different principles and objects. The main purpose of mediation analysis is to identify and estimate various average causal mediation effects (ACMEs). Identification of these effects relies on the assumption of sequential ignorability (Imai and Yamamoto, 2013). On the other hand, our framework aims to infer the activation of a mechanism (for at least one unit in the sample) by using heterogeneous treatment effects, which instead, relies on exclusion assumptions that we propose (Assumptions 1-2). The following DAGs facilitate our discussion of the differences in these approaches.
@@ Appendix D Comparison to Mediation Analysis
Consider first the case with multiple independent causal mechanisms in Figure A.2. In the left DAG, treatment $Z$ indirectly affects outcome $Y$ through two channels $M_{1}$ and $M_{2}$, and may also directly affect $Y$. To non-parametrically identify average indirect effect mediated by $M_{1}$, the key part of the sequential ignorability is $Y_{i}(z^{\prime},m_{1},M_{2i}(z^{\prime}))\perp M_{1i}|Z_{i}=z$. The assumption is challenging to interpret and cannot be guaranteed by most experimental designs because it involves cross-indices independence, from $z$ to $z^{\prime}$. Graphically, sequential ignorability requires that all variables $U$ should be observed and included in the adjustment set. Another important implication of the assumption is that $U$ cannot be affected by the treatment $Z$, i.e., the dashed line is not allowed. When these assumptions hold, mediator $M_{1}$ must be measured in order to estimate the ACME.
@@
In the right DAG, treatment $Z$ again affects outcome $Y$ through two channels $M_{1}$ and $M_{2}$, and may also directly affect $Y$. Mechanism $M_{1}$ is activated if its average indirect effect is non-zero for some unit. The existence of HTEs provides a sufficient condition for this activation under both exclusion assumptions. In our framework, variables $U$ may or may not be measured or included in the adjustment set. Further, $U$ can be a child of treatment, $Z$ (though this introduces a third mechanism). For HTE to (ever) be informative of mechanism activation, we need to observe another pre-treatment variable $X$. It is assumed not to moderate (average) direct effect and (average) indirect effect mediated by $M_{2}$. That is, two dashed lines are excluded in the right DAG in Figure A.2. From these DAGs, it is clear that there is no logical ordering of the strength of the two types of assumptions.
@@
There are many other differences between the two methods. For example, in the mediation analysis, mediators must be measurable and measured while these measurements are not required by our framework. However, in our framework, researchers must have a measured candidate MIV $X$ that is believed to satisfy Assumptions 1-2. Also, when using HTE to detect mechanisms, researchers need to pay more attention to whether $Y$ is directly affected outcome. Even though we have emphasized their differences, two frameworks also have shared features. For example, both require that the total causal effect of $Z$ on $Y$ is identified. This is explicitly assumed by sequential ignorability $\{Y_{i}(z,m_{1i},m_{2i}),M_{1i},M_{2i}\}\perp Z_{i}$ and implicitly assumed in our framework.
D.1 Related Mechanisms and Correlated Mediators
@@
Because extension to related mechanisms (correlated mediators) is not our main focus, we only make some brief comments. Consider two different correlation structures. In the left DAG of Figure A.3, mediator $M_{2}$ directly affects $M_{1}$. As mentioned in the Imai and Yamamoto (2013), two assumptions are required to identify the ACME with respect to $M_{1}$. The first one is the modified sequential ignorability assumption. Unfortunately, with causally dependent multiple mediators, an assumption of no treatment-mediator interaction effects is also required. For the HTE-mechanism framework, we can simply treat correlated mechanisms as one (molar) mechanism. Then, as long as exclusion assumptions hold for the average direct effect and other indirect effects, our results in the main text still hold. One caveat is that if the $M_{1}$ mechanism is inactive, for example, because the dashed line in the figure disappears, then the HTE-mechanism framework may yield misleading inferences about the influence of mechanism 1.
@@
Appendix E Proofs of Propositions
E.1 Proof of Proposition 1
@@
Equation (29) follows from the linearity of expectations and the decomposition of the total effect in 11. Equation (30) is guaranteed by the linearity of $L$.
Then, under exclusion assumptions 1 and 2, we can express:
@@
HTE exist with respect to $X_{k}$ if (32) is non-zero. In this case, then $X_{k}\in\textbf{X}_{j}^{MIV}$ by the definition of MIV.
E.2 Proof of Propostions 2
@@
Prove by contrapositive. Suppose not, which means $\textbf{X}_{j}^{MIV}$ is non-empty and for some $x,x^{\prime}\in\mathbb{R}$, $IE_{j}(X_{k}=x)\neq IE_{j}(X_{k}=x^{\prime})$. Then:
@@
Note $E_{X_{\neg k}}[DE(X_{k}=x^{\prime})]=ADE(X_{k}=x)$ and $E_{X_{\neg k}}[IE_{j}(z,z^{\prime};X_{k}=x)]=AIE_{j}(X_{k}=x^{\prime})$. Equation (35) follows because $ADE(X_{k}=x)=ADE(X_{k}=x^{\prime})$ and $AIE_{i\neq j}(X_{k}=x)=AIE_{i\neq j}(X_{k}=x^{\prime})$ under exclusion assumptions 1 and 2. We find HTE with respect to $X_{k}$.
So, we have shown that if two conditions do not hold, then HTE exists for $X_{k}=x$ and $X_{k}=x^{\prime}$. By contrapositive, we prove that if no HTEs exist with respect to $X_{k}$, at least one of the two conditions must be true.
E.3 Proof of Proposition 3
@@
By definition of $\textbf{X}^{R}$, we know $X\notin\textbf{X}^{R}$ implies that $X$ must be independent of $Y$. We prove this proposition by contrapositive. Let $\mathbb{P}(\tilde{y}|Z,X_{k})$ be the conditional distribution of $h(Y)$. Suppose $X_{k}\notin\textbf{X}^{R}$, then:
@@
Given pre-treatment variables {$_{1},X_{2},...,X_{n}\}=\textbf{X}$ X and outcome variable $Y$. Variables {$_{1},...,X_{m}\}\subset\textbf{X}$ X are MIVs for mechanism j (denote them as the set $\textbf{X}_{j}^{MIV}$ and the remaining as the set $\textbf{X}_{j}^{non-MIV}$) if and only if there exits function $g_{1}(\cdot)$ and non-additively separable function $g_{2}(\cdot)$, and $Y$ satisfies:
$Y=g_{1}(\textbf{X}_{j}^{non-MIV},\textbf{X}_{j}^{MIV})+g_{2}(\textbf{X}_{j}^{MIV},Z)$ (41)
@@
    The non-additively separable function $g_{2}(\textbf{X}_{j}^{MIV},Z)$ can take the form $g_{3}(T)+g_{4}\textbf{X}_{j}^{MIV},T)$ for some function $g_{3}(\cdot)$ and non-additively separable function $g_{4}(\cdot)$.
@@
$CATE(X=x)-CATE(X=x^{\prime})=E[h_{2}(\textbf{X}_{j}^{MIV},Z)|X=x]-E[h_{2}(\textbf{X}_{j}^{MIV},Z^{\prime})|X=x^{\prime}]$ (42)
Equation (42) is only a function $\textbf{X}_{j}^{MIV}$ because $h_{1}(\cdot)$ cancels out.
@@
E.4 Extension of Proposition 4
In the main text, Proposition 4 indicates that if there exist no HTE for the indirectly affect outcome, $X_{k}$ can be any relevant or non-relevant covariate. Now we provide a stronger version of Proposition 4 by imposing assumptions about the directly-affected outcome, $Y$, and the form of the non-linear transformation $h(Y)$. These assumptions permit additional learning from the lack of HTE in this case.
In practice, most indirectly affected outcomes are discrete variables, such as voting behavior, survey responses, or choices. Let us consider the following non-linear transformation of the directed affected outcome $Y$:
$h(Y)=\begin{cases}y_{1}&Y\in(-\infty,c_{1}]\\ y_{2}&Y\in(c_{1},c_{2}]\\ ...\\ y_{q}&Y\in(c_{q+1},\infty)\\ \end{cases}$ (43)
@@ E.4 Extension of Proposition 4
If $X_{k}$ is not effective, then $CATE(X_{2}=x)=CATE(X_{2}=x^{\prime})$, therefore there exist no HTE. If $X_{k}$ is effective, then non-existence of HTE requires that
@@
Appendix F Strengthening Assumptions
@@
Recall the basic problem with indirectly-affected outcomes. In Figure A.4, $Y$ is the directly-affected outcome, and $h(Y)$ is the non-linear transformation of the directly-affected outcome. Our main result shows that $X^{non-MIV}$ can also induce HTE even though $X^{non-MIV}$ does not indicate the mechanism.
@@
Figure A.4: Indirectly-affected outcome DGP.
F.1 Discrete Outcomes under Monotonicity Assumptions
@@
[…]
Figure A.5: Two different DGPs. On the left, in DGP 1, $X$ is a MIV. On the right, in DGP 2, $X$ is not a MIV.
The left panel assumes $X$ is a MIV. $X$ is not a MIV in the right panel. In the figure, there are no other mediators. Therefore, both graphs satisfy exclusion assumptions 1 and 2 by construction.
We will assume that $Y_{s}$ is a latent directly-affected outcome and $h(Y_{s})$ is the observed binary variable:
@@
for some $c\in(-\infty,\infty)$. Propositions 3 and 4 show that we cannot differentiate between the left and right on the basis of the existence or non-existence of HTE for indirectly-affected outcome $h(Y_{s})$.
We will consider what can be gained by imposing a monotonicity assumption of the form: $\frac{\partial^{2}Y}{\partial Z\partial X}>(<)0$ (note that the inequalities are strict). [^21] Writing the monotonicity assumption in this way assumes that this derivative exists. Clearly montonicity can hold in the left panel (where $X$ is a MIV) but $\frac{\partial^{2}Y}{\partial Z\partial X}=0$ in the right panel in which $X$ is not a MIV. To explore the implications of monotonicity, we consider following DGPs for $i=\{1,2\}$:
@@
For $X\in X^{MIV}$, e.g., the left DGP in Figure A.5, suppose that montonicity holds such that $\frac{\partial^{2}Y_{1}}{\partial X\partial Z}:=\beta(x,z)$, where $\beta(x,z)$ is either strictly positive or negative. For $X\notin X^{MIV}$, e.g., the right DGP in Figure A.5, by definition, we have $\frac{\partial^{2}Y_{2}}{\partial x\partial z}=0$.
@@
Let $f_{e}$ be the density of $e$ denote its derivative as $f^{\prime}_{e}$. We can then express HTE for $h(Y_{1})$ as:
@@
(1) Suppose that $e$ is uniformly distributed, then HTE for $h(Y_{2})$ is 0.
(2) Suppose $e$ is not uniformly distributed, then HTE for $h(Y_{1})$ and $h(Y_{2})$ have different signs under two cases:
@@
F.2 Assumptions on the Latent Utility Distributions
In practice, one of the most common cases is that $Y$ is the latent utility, and $h(Y)$ is the observed action or discrete choice. If we can re-construct the utility from the observed data, then we can use it as the directly-affected outcome and use HTE to assess mechanism activation under Assumptions 1-2. From Propositions 1 and 2, we know more information on mechanism activation can be ascertained from HTE on directly-affected outcome.

### v2

@@ Appendix D Directly/Indirectly-Affected versus Latent/Observed Outcomes
Appendix E Comparison to Mediation Analysis
In this section, we compare our framework connecting HTEs and mechanisms to mediation analysis. It is important to note that the two frameworks are built on different principles and objects. The main purpose of mediation analysis is to identify and estimate various average causal mediation effects (ACMEs). Identification of these effects relies on the assumption of sequential ignorability (Imai and Yamamoto, 2013). On the other hand, our framework aims to infer the activation of a mechanism (for at least one unit in the sample) by using heterogeneous treatment effects, which instead, relies on exclusion assumptions that we propose (Assumptions 2-3). The following DAGs facilitate our discussion of the differences in these approaches.
@@ Appendix E Comparison to Mediation Analysis
Consider first the case with multiple independent causal mechanisms in Figure A.3. In the left DAG, treatment $Z$ indirectly affects outcome $Y$ through two channels $M_{1}$ and $M_{2}$, and may also directly affect $Y$. To non-parametrically identify average indirect effect mediated by $M_{1}$, the key part of the sequential ignorability is $Y_{i}(z^{\prime},m_{1},M_{2i}(z^{\prime}))\perp M_{1i}|Z_{i}=z$. The assumption is challenging to interpret and cannot be guaranteed by most experimental designs because it involves cross-index independence, from $z$ to $z^{\prime}$. Graphically, sequential ignorability requires that all variables $U$ should be observed and included in the adjustment set. Another important implication of the assumption is that $U$ cannot be affected by the treatment $Z$, i.e., the dashed line is not allowed. When these assumptions hold, mediator $M_{1}$ must be measured in order to estimate the ACME.
@@
In the right DAG, treatment $Z$ again affects outcome $Y$ through two channels $M_{1}$ and $M_{2}$, and may also directly affect $Y$. Mechanism $M_{1}$ is activated if its average indirect effect is non-zero for some unit. The existence of HTEs provides a sufficient condition for this activation under both exclusion assumptions. In our framework, variables $U$ may or may not be measured or included in the adjustment set. Further, $U$ can be a child of treatment, $Z$ (though this introduces a third mechanism). For HTEs to (ever) be informative of mechanism activation, we need to observe another pre-treatment variable $X$. It is assumed not to moderate (average) direct effect and (average) indirect effect mediated by $M_{2}$. That is, two dashed lines are excluded in the right DAG in Figure A.3. From these DAGs, it is clear that there is no logical ordering of the strength of the two types of assumptions.
@@
There are many other differences between the two methods. For example, in the mediation analysis, mediators must be measurable and measured while these measurements are not required by our framework. However, in our framework, researchers must have a measured candidate MDV $X$ that is believed to satisfy Assumptions 2-3. Also, when using HTEs to detect mechanisms, researchers need to pay more attention to whether $Y$ is directly affected outcome. Even though we have emphasized their differences, two frameworks also have shared features. For example, both require that the total causal effect of $Z$ on $Y$ is identified. This is explicitly assumed by sequential ignorability $\{Y_{i}(z,m_{1i},m_{2i}),M_{1i},M_{2i}\}\perp Z_{i}$ and implicitly assumed in our framework.
E.1 Related Mechanisms and Correlated Mediators
@@
Because extension to related mechanisms (correlated mediators) is not our main focus, we only make some brief comments. Consider two different correlation structures. In the left DAG of Figure A.4, mediator $M_{2}$ directly affects $M_{1}$. As mentioned in the Imai and Yamamoto (2013), two assumptions are required to identify the ACME with respect to $M_{1}$. The first one is the modified sequential ignorability assumption. Unfortunately, with causally dependent multiple mediators, an assumption of no treatment-mediator interaction effects is also required. For the HTE-mechanism framework, we can simply treat correlated mechanisms as one (molar) mechanism. Then, as long as exclusion assumptions hold for the average direct effect and other indirect effects, our results in the main text still hold. One caveat is that if the mechanism represented by $M_{1}$ is inactive, for example, because the dashed line in the figure disappears, then the HTE-mechanism framework may yield misleading inferences about the influence of mechanism 1.
@@
Appendix F Proofs of Propositions
We prove propositions under general multi-mechanism cases, as stated in the C. All propositions in the main text are all re-stated in the special case of one mechanism. Assumption 3 is only invoked in the multiple mechanism generalization.
F.1 Proof of Proposition 1
Proposition 5.
Suppose that $Y$ is directly affected by mechanism $j$ and Assumptions 2 and 3 hold with respect to $X_{k}$. If HTEs exist with respect to $X_{k}$, then $X_{k}\in\textbf{X}_{j}^{MDV}$ for mechanism $j$.
@@ F.1 Proof of Proposition 1
Equation (36) follows from the linearity of expectations and the decomposition of the total effect in 31. Equation (37) is guaranteed by the linearity of $L$.
Then, under exclusion assumptions 2 and 3, we can express:
@@
HTEs exist with respect to $X_{k}$ if (39) is non-zero. In this case, then $X_{k}\in\textbf{X}_{j}^{MDV}$ by the definition of MDV.
F.2 Proof of Propostions 2
Proposition 6.
  1. 1.
    $X_{k}\notin\textbf{X}^{MDV}$ for mechanism $j$.
  2. 2.
    No MDV exists.
@@ F.2 Proof of Propostions 2
Proof by contrapositive. Suppose not, which means $\textbf{X}_{j}^{MDV}$ is non-empty and for some $x,x^{\prime}\in\mathbb{R}$, $IE_{j}(X_{k}=x)\neq IE_{j}(X_{k}=x^{\prime})$. Then:
@@
Note $E_{X_{\neg k}}[DE(X_{k}=x^{\prime})]=ADE(X_{k}=x)$ and $E_{X_{\neg k}}[IE_{j}(z,z^{\prime};X_{k}=x)]=AIE_{j}(X_{k}=x^{\prime})$. Equation (42) follows because $ADE(X_{k}=x)=ADE(X_{k}=x^{\prime})$ and $AIE_{i\neq j}(X_{k}=x)=AIE_{i\neq j}(X_{k}=x^{\prime})$ under exclusion assumptions 2 and 3. We find HTEs with respect to $X_{k}$.
We have thus shown that if two conditions do not hold, then HTEs exist for $X_{k}=x$ and $X_{k}=x^{\prime}$. By contrapositive, we prove that if no HTEs exist with respect to $X_{k}$, at least one of the two conditions must be true.
F.3 Proof of Proposition 3
Proposition 7.
@@ F.3 Proof of Proposition 3
Proof by contrapositive. By definition of $\textbf{X}^{R}$, we know $X\notin\textbf{X}^{R}$ implies that $X$ must be independent of $Y$. Let $\mathbb{P}(\tilde{y}|Z,X_{k})$ be the conditional distribution of $h(Y)$. Suppose $X_{k}\notin\textbf{X}^{R}$, then:
@@
Given pre-treatment variables {$_{1},X_{2},...,X_{n}\}=\textbf{X}$ X and outcome variable $Y$. Variables {$_{1},...,X_{m}\}\subset\textbf{X}$ X are MDVs for mechanism j (denote them as the set $\textbf{X}_{j}^{MDV}$ and the remaining as the set $\textbf{X}_{j}^{non-MDV}$) if and only if there exits function $g_{1}(\cdot)$ and non-additively separable function $g_{2}(\cdot)$, and $Y$ satisfies:
$Y=g_{1}(\textbf{X}_{j}^{non-MDV},\textbf{X}_{j}^{MDV})+g_{2}(\textbf{X}_{j}^{MDV},Z)$ (48)
@@
    The non-additively separable function $g_{2}(\textbf{X}_{j}^{MDV},Z)$ can take the form $g_{3}(T)+g_{4}(\textbf{X}_{j}^{MDV},T)$ for some function $g_{3}(\cdot)$ and non-additively separable function $g_{4}(\cdot)$.
@@
$CATE(X=x)-CATE(X=x^{\prime})=E[h_{2}(\textbf{X}_{j}^{MDV},Z)|X=x]-E[h_{2}(\textbf{X}_{j}^{MDV},Z^{\prime})|X=x^{\prime}]$ (49)
Equation (49) is only a function of $\textbf{X}_{j}^{MDV}$ because $h_{1}(\cdot)$ cancels out.
@@
F.4 Extension of Proposition 4
Proposition 8.
Proposition 8 indicates that if there exist no HTEs for the indirectly-affected outcome, $X_{k}$ can be any relevant or non-relevant covariate. Now we provide a stronger version of Proposition 8 by imposing assumptions about the directly-affected outcome, $Y$, and the form of the non-linear transformation $h(Y)$. These assumptions permit additional learning from the lack of HTEs in this case.
In practice, most indirectly-affected outcomes are discrete variables, such as voting behavior, survey responses, or choices. Let us consider the following non-linear transformation of the directed affected outcome $Y$:
$h(Y)=\begin{cases}y_{1}&Y\in(-\infty,c_{1}]\\ y_{2}&Y\in(c_{1},c_{2}]\\ ...\\ y_{q}&Y\in(c_{q-1},\infty)\\ \end{cases}$ (50)
@@ F.4 Extension of Proposition 4
If $X_{k}$ is not effective, then $CATE(X_{2}=x)=CATE(X_{2}=x^{\prime})$, therefore there exist no HTE. If $X_{k}$ is effective, then non-existence of HTEs requires that
@@
Appendix G Strengthening Assumptions
@@
Recall the basic problem with indirectly-affected outcomes. In Figure A.5, $Y$ is the directly-affected outcome, and $h(Y)$ is the non-linear transformation of the directly-affected outcome. Our main result shows that $X^{k+1}$ can also induce HTEs even though $X^{k+1}$ does not detect the mechanism.
@@
Figure A.5: DGP for an indirectly-affected outcome, $h(Y)$.
G.1 Discrete Outcomes under Monotonicity Assumptions
@@
$Z$$M$$Y_{1}$$h(Y_{1})$$X$$Z$$M$$Y_{2}$$h(Y_{2})$$X$
Figure A.6: Two different DGPs. On the left, in DGP 1, $X$ is a MDV. On the right, in DGP 2, $X$ is not a MDV.
The left panel assumes $X_{k}$ is a MDV. $X_{k+1}$ is not a MDV in the right panel. In the figure, there are no other mediators. Therefore, both graphs satisfy exclusion assumptions 1 and 2 by construction.
We will assume that $Y_{s}$ is a latent directly-affected outcome and $h(Y_{s})$ is a binary variable given by:
@@
for some $c\in(-\infty,\infty)$. Propositions 3 and 4 show that we cannot differentiate between the DGPs 1 and 2 in Figure A.6 on the basis of the existence or non-existence of HTEs for indirectly-affected outcome $h(Y_{s})$.
We will consider what can be gained by imposing a monotonicity assumption of the form: $\frac{\partial^{2}Y}{\partial Z\partial X}>(<)0$ (note that the inequalities are strict). [^18] Writing the monotonicity assumption in this way assumes that this derivative exists. Clearly montonicity can hold in the left panel (where $X$ is a MDV) but $\frac{\partial^{2}Y}{\partial Z\partial X}=0$ in the right panel in which $X$ is not a MDV. To explore the implications of monotonicity, we consider following DGPs for $s=\{1,2\}$:
@@
For $X\in X^{MDV}$, e.g., the DGP 1 in the left panel of Figure A.6, suppose that montonicity holds such that $\frac{\partial^{2}Y_{1}}{\partial X\partial Z}:=\beta(x,z)$, where $\beta(x,z)$ is either strictly positive or negative. For $X\notin X^{MDV}$, e.g., the right DGP in Figure A.6, by definition, we have $\frac{\partial^{2}Y_{2}}{\partial x\partial z}=0$.
@@
Let $f_{e}$ be the density of $e$ denote its derivative as $f^{\prime}_{e}$. We can then express HTEs for $h(Y_{1})$ as:
@@
(1) Suppose that $e$ is uniformly distributed, then HTEs for $h(Y_{2})$ is 0.
(2) Suppose $e$ is not uniformly distributed, then HTEs for $h(Y_{1})$ and $h(Y_{2})$ have different signs under two cases:
@@
Proposition 10 reveals and Corollary A1 illustrates that Assumptions 2-3 and monotonicity are not, in general, jointly sufficient to use HTEs to assess mechanism activation on indirectly-affected outcomes.
G.2 Assumptions on the Latent Utility Distributions
In practice, one of the most common cases is that $Y$ is the latent utility, and $h(Y)$ is the observed action or discrete choice. If we can re-construct the utility from the observed data, then we can use it as the directly-affected outcome and use HTEs to assess mechanism activation under Assumptions 1-2. From Propositions 1 and 2, we know more information on mechanism activation can be ascertained from HTEs on directly-affected outcome.